OpenStudy (anonymous):

4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

4(x − 1) = 2(x + 2) + 6 4x-4=2x+4+6

OpenStudy (anonymous):

Now, can u?

OpenStudy (kaederfds):

i got x=3

OpenStudy (anonymous):

Nope..... @Kaederfds

OpenStudy (anonymous):

4x-4=2x+4+6 4x-2x=4+6+4 2x=14

OpenStudy (monica0597):

start by using the distributive property.